OpenStudy (anonymous):

For any constant a, let g(x) = ax − 2xln(x) for x > 0. (a) What is the x-intercept g(x)? x =

OpenStudy (anonymous):

solve \[ax+2x\ln(x)=0\] i guess

OpenStudy (anonymous):

\[2x\ln(x)=-ax\] is a start then since you know \(x\neq 0\) i think you can go ahead and divide both sides by \(x\) and get \[2\ln(x)=-a\] or \[\ln(x)=-\frac{a}{2}\]

OpenStudy (anonymous):

hmm maybe that is wrong maybe factoring is better \[x(a+2\ln(x))=0\] nope works the same way

OpenStudy (anonymous):

yeah I had that as my original answer and got it wrong ;/ I'm lost on how to do it a different way

OpenStudy (anonymous):

Im also trying to find the x coordinate of the critical point of g(x)

OpenStudy (freckles):

to find the x-intercept set y=0 and solve for x

OpenStudy (freckles):

\[ax-2x \ln(x)=0\\ x(a-2 \ln(x))=0 \\ x \neq 0 \\ a-2 \ln(x)=0 \\ \text{ finish solving for x }\]

OpenStudy (freckles):

\[2 \ln(x)=a \]

OpenStudy (anonymous):

so e^a/2?

OpenStudy (freckles):

yeah (e^(a/2), 0) <--x-intercept

OpenStudy (anonymous):

okay thanks! so to find the x vale of the critical point of g(x) i just set the first derivative to zero and solve for x..?

OpenStudy (freckles):

solve g'=0 g' dne and make sure those numbers occur in the domain those are critical numbers

OpenStudy (anonymous):

ahhh i see! Okay thank you! I was able to solve it!

OpenStudy (freckles):

coolness